# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### 12 Mar 2026 v3

@@ Abstract
We present the exact analytic solution of the Makeenko-Migdal loop equations, thereby solving planar QCD $(N_{c}\rightarrow\infty)$ in the continuum limit. The solution is constructed by quantizing internal Majorana fermions (elves) on a rigid Hodge-dual minimal surface. We demonstrate that these worldsheet degrees of freedom provide the exact algebraic mechanism required to satisfy the unintegrated vector loop equations, with the Pauli principle enforcing the planar factorization. In the local limit, the theory reduces to a confining analytic twistor string. By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is exactly governed by Catastrophe Theory and classified by the topological number of twistor poles. The simplest sector with just one pole yields parametric Regge trajectories governed by trigonometric functions. While mathematically non-linear, these geometric trajectories are numerically remarkably close to linear, matching the experimental data for light mesons (the unshifted $\pi,K$ trajectories and the topologically shifted $\rho$ trajectory) within 95% confidence intervals. The empirically correct intercepts $\alpha_{\pi}(0)=-1/96$ and $\alpha_{\rho}(0)=+47/96$ emerge directly and purely algebraically from the 2D Liouville kinetic anomaly interacting with the twistor branch point monodromy, without any ad hoc assumptions of string vibrations. Ultimately, this exact geometric quantization explicitly realizes the Master Field as a critical classical trajectory in twistor space.
@@ 1 Introduction
The identification of the string description of large-$N_{c}$ QCD has been a central challenge of strong-interaction theory since the advent of the MM loop equations [^1] The Makeenko-Migdal equations, referred to hereafter as the MM equations, exhibit the characteristic factorization properties of a string theory. While brevity dictates this acronym, the foundational credit remains fully shared with my former student, Yuri Makeenko, whose hard work and physical intuition half a century ago made these derivations possible.. A concrete, mathematically controlled continuum realization has remained elusive.
Historically, the main obstruction is the ultraviolet singular nature of the coordinate-space Wilson loop functional $W[C]$. Even after renormalization, $W[C]$ exhibits cusp and perimeter divergences, and the loop equation contains a contact term $\delta^{(4)}(x-y)$ at self-intersections. In the standard coordinate representation, this makes the local limit technically delicate and conceptually opaque: the loop equation is local, while the perturbative solution is nonlocal and develops logarithmic singularities tied to cusp geometry.
In this paper, we adopt a different viewpoint: rather than attempting to “repair” the coordinate-space Wilson loop as an observable object, we formulate the dynamics directly in momentum loop space. The momentum loop functional $W[P]$ was introduced in earlier work as the functional Fourier transform of $W[C]$. Its decisive advantage is structural: the planar loop equation in momentum space becomes an exact algebraic-differential functional equation with no coordinate-space delta functions. In this representation, the geometric cusp singularities of $W[C]$ are integrated out, and the continuum equations admit a local formulation that is finite in the sense relevant for physical quark-loop amplitudes.
The goal of this second paper in the Geometric QCD series is to provide the dynamics that complement the kinematics established in Part I [36]. Part I constructs the rigid vacuum geometry: the Hodge-dual additive minimal surface $S_{\chi}[C]$ whose (anti)self-dual area derivative yields a zero mode of the loop operator, producing the confining dressing factor $\exp\left(-\kappa S[C]\right)$. Here, we attach dynamical worldsheet degrees of freedom to this rigid background without summing over fluctuating metrics and show how the (regularized) planar loop equations are satisfied by this Ansatz for $W[C]$.
Our framework synthesizes three ingredients:
@@
    Fermi string (“Elfin theory”). We revisit the fermionic construction in which internal Majorana degrees of freedom on the worldsheet implement the required sign structure at intersections. In particular, the Pauli principle enforces cancellations between nonplanar intersection configurations, leaving the planar factorization pattern.
@@
    Rigid Hodge-dual minimal surface. The fermions are not placed on a random surface. Instead, they live on the rigid Hodge-dual minimal surface determined holographically by the boundary loop. This removes the dynamical Liouville instability that afflicted the original random-surface formulation: there is no summation over worldsheet metrics, and the geometry is fixed by the loop data.
@@
    Momentum loop dynamics and the twistor resolution. In momentum loop space, the loop equation becomes an exact algebraic-differential equation, stripped of coordinate-space contact singularities. We rigorously demonstrate that while this equation admits a functional Taylor-Magnus expansion at low orders, a purely one-dimensional algebraic Master Field solution violently breaks down at the eighth order ($\mathcal{W}^{(8)}$). This mathematical catastrophe proves that a 1D topological string fundamentally lacks the internal degrees of freedom to absorb the macroscopic spatial stress of Planar QCD. To resolve this geometric anomaly, the reduced loop-space measure must be parametrized by 4D continuous boundary twistor variables. This leads to an effective “analytic twistor string” description: a boundary sigma-model (a coset built from two $S^{3}$ spinors modulo a local $U(1)$) coupled to a holographic Liouville field whose bulk values are determined by the unique holomorphic extension of the boundary twistors.
@@
Equation solving vs. model building The starting point of this work is not a proposed “string model of QCD,” but the planar MM loop equations, treated as the nonperturbative defining equations of large-$N_{c}$ Yang–Mills/QCD. The aim is constructive: to exhibit explicit variables and an explicit functional measure in which the loop equations admit a finite continuum solution and yield physical observables. The principal ingredients of the present construction are not adjustable assumptions. The rigid Hodge-dual minimal surface enters because strict additivity and an (anti)self-dual area derivative are algebraically required for compatibility with the loop hierarchy, while the Majorana determinant enters because its Pauli-principle sign structure enforces planar factorization.
[…]
The culminating result of this paper is the exact reduction of the meson spectroscopy problem to a one-dimensional functional integral over boundary twistor data $(\lambda(z),\mu(z),|z|=1)$. By analyzing the monodromies of the complexified effective action, we reveal that the discrete mass spectrum is deterministically governed by Catastrophe Theory. The path integral exactly localizes at critical resonance energies where the Lefschetz thimbles become degenerate. Because the integration variables are coupled holomorphic twistors, this degeneracy intrinsically possesses a corank of 2, mathematically upgrading generic semiclassical branch cuts into the exact simple poles of hadronic states. Planar QCD thus ceases to be a theory of stochastic quantum fluctuations; the extraction of the exact spectrum reduces to a deterministic, classical generalized eigenvalue problem.
The $S$-matrix poles are generated by non-compact flat valleys in the effective action, which lead to the divergence of the steepest descent integral. We demonstrate that these flat valleys are dynamically produced by the multi-valued monodromy of the holomorphic twistor fields; along these winding trajectories, higher-order functional corrections to the action are strictly absent. Furthermore, the discrete spectrum is classified by the topological number of twistor poles inside the unit circle. This topological sector is rigorously protected: the crossing of the unit circle by a twistor pole leads to a divergent action, creating an impenetrable topological barrier, because the conjugate poles $z_{0}$ and $1/\bar{z}_{0}$ inevitably pinch the physical integration loop $|z|=1$. By analytically continuing the action to the boundary and summing over the infinite number of windings, with $SU(2)$ rotated boundary conditions for twistors, after projecting on the fixed spin state we obtain an exact Bohr-Sommerfeld quantization condition. For the simplest sector with just one pole, this yields the unified parametric Regge trajectory:
$J=\frac{\sin^{4}(\pi a)}{2\pi^{2}a(1-a)}\frac{m^{2}}{\sigma}-\frac{q}{2}-\frac{a}{48};$ (1.1)
$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+24(2J+q)}$ (1.2)
The topological trace phase $q=0$ identifies the unshifted unnatural parity ($\pi$, $K$) trajectories, while q=1q=-1 rigidly shifts the natural parity ($\rho$) trajectory by $1/2$.
Remarkably, while mathematically non-linear, these trajectories are numerically very close to straight lines (see Figure 15). The empirically correct asymptotic intercepts $\alpha_{\pi}(0)=-1/96$ and $\alpha_{\rho}(0)=47/96$ emerge directly from the 2D Liouville kinetic anomaly. The asymptotic expansion of the Regge trajectory $J(s)$ at $s=m^{2}/\sigma>1$ reads:
$J(s)$ $=\frac{2s}{\pi^{2}}-\left(\frac{q}{2}+\frac{1}{96}\right)+\frac{B_{1}}{s}+\frac{B_{3}}{s^{3}}+\dots);$ (1.3)
$B_{1}$ $=\frac{\pi^{2}}{36864\left(\pi^{2}-2\right)}\approx 0.0000340208;$ (1.4)
$B_{3}$ $=\frac{\pi^{6}\left(48-24\pi^{2}+5\pi^{4}\right)}{32614907904\left(\pi^{2}-2\right)^{4}}\approx 2.29*10^{-9}$ (1.5)
We interpret this construction as a geometric realization of the large-$N_{c}$ Master Field [54]. While the Master Field does not manifest as a single classical gauge configuration in coordinate spacetime, it admits an explicit formulation in loop space. The true classical limit of planar QCD is thus identified not as a matrix algebra, but as a rigid classical geometric object: a critical trajectory in twistor space, conceptually parallel to the spectral curves of Seiberg-Witten theory[45]. This geometric structure is accessible via exact WKB analysis [50, 12], where twistor monodromies dynamically generate flat valleys within the Lefschetz thimbles [55, 3].
Organization of the paper and navigation points.
To assist the reader in navigating the conceptual shifts and technical machinery of this framework, we have integrated our structural outline with five specific navigation points that address anticipated physical and computational questions.
Section 2 summarizes the loop-equation zero modes and recalls why the Hodge-dual minimal surface provides the appropriate rigid geometric background. For readers concerned about (i) the physical meaning of loop-equation zero modes, Sec. 2.3 explicitly details the underlying multi-instanton and vacuum ambiguity in loop space. Sections 3 to  9 then develop the fermionic determinant on this rigid surface and derive the corresponding loop equation, including its local limit.
Sections 10 to 12 introduce momentum loop space and the twistor string framework. To clarify (ii) the status of Wilson-loop “renormalization,” Sec. 10.1 sharply outlines our perspective: $W[C]$ is kept regularized at a finite Elf mass, physical observables are constructed natively in momentum loop space $W[P]$ (where coordinate-space singularities are integrated out), and only then is the local limit taken. Following this, Sections 11 to  12 execute the central change of variables. Addressing (iii) the parametrization of the quark loop in phase space by boundary values of holomorphic twistors, Sec. 11 details how the linear phase space measure is transformed into twistor variables by factoring out the volume of the diffeomorphism group using the Virasoro constraint.
Section 13.1 presents the twistor-holographic (“gauge holography”) interpretation and clarifies the relation of the resulting analytic confining twistor string to standard holography and to topological twistor strings.
Finally, we highlight (iv) the exact analytic extraction of the mass spectrum. Section 15 reformulates the whole problem of the QCD mass spectrum as a complex saddle point equation with a degenerate Hessian, corresponding to the collision of two saddle points in the spirit of Catastrophe theory. Section 16 then reframes this localization as an infinite-dimensional realization of Picard-Lefschetz theory, Resurgence, and Exact WKB analysis. We establish a strict structural correspondence between the twistor monodromies of Geometric QCD and the spectral networks of Seiberg-Witten theory, identifying the exact resonance condition as the twistor-string analogue of an Argyres-Douglas point. Section 17 concludes by synthesizing these results into the final geometric Master Field. The appendices collect vital technical material, including the exact algebraic demonstration of the $\mathcal{W}^{(8)}$ Master Field breakdown in C and computations of the Minkowski action for our chiral helicoid minimal surface.
@@
Conceptual Demarcation: Bypassing 2D Quantum Gravity
Because the mathematical vocabulary of this paper (strings, twistors, minimal surfaces) intersects with decades of literature on string theory, a reader may instinctively attempt to map this construction onto the conventional quantization of a fluctuating 2D worldsheet. To prevent fundamental conceptual misunderstandings, it is vital to explicitly state that we are not proposing a phenomenological string model, nor are we quantizing a fundamental string in the Polyakov sense.Consequently, the reader will not find integrations over fluctuating 2D bulk metrics, conformal ghost sectors, or integrations over ambiguous Teichmüller moduli spaces. Attempting to evaluate this work by the standard criteria of abstract 2D string modeling is a category error.Our absolute starting point is the rigorously defined, 1-dimensional phase space of Planar QCD: the momentum-loop representation of the MM equations, governed by the exact linear measure in phase space $\int\mathcal{D}C^{\prime}\mathcal{D}P$. The transition to "twistor space" is strictly a local, 1-dimensional algebraic change of variables on the boundary circle. The twistors do not dynamically fluctuate across a 2D bulk; they are localized strictly at the 1D boundary to parametrize the loop velocity, and are analytically continued inward solely as a tool of complex analysis. We are transforming the intractable functional geometry of a 1D gauge-theory phase space into a rigorously solvable problem of 1D complex analysis and linear algebra.
@@ 2 Zero modes of Loop equation
Let us summarize the theory of Hodge-dual minimal surfaces and its relevance to the MM loop equation. For a detailed discussion and proofs of relevant theorems, the reader may look into the latest paper [36], and the details of the new loop calculus without divergences can be found in [33, 37].
@@ 2.1 Self-dual area derivative and loop equation
Let us recall why we are looking for the self-dual minimal surface, as suggested in [33]. In this section, we follow the arguments of our recent paper [36], but our goal is not to present the Hodge-dual surface per se, but rather to view it as a hologram of the QCD string, with Fermi degrees of freedom attached.
The area derivative for a general functional $W[C]$ was defined in that paper as a discontinuity of the second variation of the area
@@
This definition implements the geometric definition of [24, 27] and is formally equivalent to the functional derivative implementation by Polyakov [41]. The crucial difference is the lack of divergences in this new definition and its unbroken parametric invariance. In the old, singular implementations, the zero modes were obscured by kinematic divergencies, interfering with potential small-distance singularities of perturbative QCD.
The new loop equation for the Yang–Mills gradient flow was shown in [33] to reduce to the loop space diffusion equation, which allows the solution $W=\exp\left(-\kappa S[C]\right)$ provided that $S[C]$ is a zero mode of this equation
@@
The basic observation made in that paper was that there is a Hodge-dual solution similar to the Yang–Mills multi-instanton. The loop space diffusion equation (2.3) is identically satisfied for any functional $S[C]$ with self-dual (SD) or anti-dual (ASD) area derivative
@@
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [33] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents the SD or ASD tensor. This choice would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
The existence of this zero mode is based on the so-called Leibniz property of the loop operator and similar singular loop operators in loop calculus.
@@
It was heuristically claimed in the original MM papers [25, 27], and rigorously proven in the loop calculus [33], section 3, eqs (39), (40). Due to the Leibnitz property,
@@
We are looking for such a self-dual minimal surface which is additive at the self-intersecting loops. This condition will be discussed in detail below. It is critical for the whole theory—due to this additivity, the confining factor $\exp\left(-\kappa S[C]\right)$ is compatible with the whole chain of the loop equations.
2.2 Compatibility with the MM loop equations.
The full chain of loop equations for the Wilson loops with a finite number of colors $N_{c}$ [25, 27] has the same loop operator (2.3) as the Yang–Mills gradient flow, except there is no time derivative. In this paper, we are only considering the leading order of the $1/N_{c}$ expansion. We use the loop calculus developed in the paper [33], as it applies to the loop diffusion operator (2.3), which is the same in the multiloop MM chain and in the Yang–Mills gradient flow equation. The MM loop equation has the structure (see Fig.1)
Figure 1: The MM equation, with the loop diffusion operator on the left side, and the integral term on the right side. The double line on the right stands for the delta function $\delta^{4}(x-y)$ enforcing self-intersection.
@@
Here $\lambda=N_{c}g_{0}^{2}$ is the ’t Hooft’s coupling constant in the limit of $N_{c}\to\infty$. This equation is singular at self-intersections, so it must be understood in the sense of distributions in loop space. It is manifestly parameterization-invariant, and the confining factor $\exp\left(-\kappa S[C]\right)$ factors out from the right side by additivity of $S[C]$ and cancels with the same factor on the left side.
Let us stress that this factorization takes place independently of the splitting points $x,y\in C$. For every self-intersection $x,y\in C$ such that $C(x)=C(y)$, the product of two factors $\exp\left(-\kappa S[C_{xy}]\right)$ and $\exp\left(-\kappa S[C_{yx}]\right)$ exactly equals the factor $\exp\left(-\kappa S[C]\right)$ on the left side of the MM equation; Thus, it can be taken out of the integral, after which it cancels on both sides of the MM equation.
The physical hypothesis of that paper was that the actual gluonic vacuum of QCD is described by this dressed solution, choosing the "fluctuative" factor $W_{\text{fluct}}$ as the seed. While the symmetry is exact, the separation into a fluctuative factor and a purely geometric area law is ambiguous. It is physically most distinct in the large-loop limit, where the area term dominates over $W_{\text{fluct}}$, which is assumed only to contribute the term proportional to a perimeter of the loop.
In the present paper, we construct an explicit solution for this factor $W_{\text{fluct}}$, adapting the fermionic determinant from the ’81 paper [28] to the Hodge-dual minimal surface instead of a random surface, assumed in that work. The confining factor becomes the vacuum energy of that rigid Fermionic string, and planar graphs emerge as a strong coupling expansion of that 2D fermi system in terms of a sum over fermionic loops on a Hodge-dual minimal surface. The Pauli principle leads to repulsion of these loops, making them topologically equivalent to planar graphs of QCD.
The comparison of the contour equation for these fermionic determinants from the ’81 work with the MM equation establishes the correspondence between the ’t Hooft coupling constant $\lambda$ and the parameters of the Elfin theory (the bare fermionic mass and the vacuum energy).
2.3 The Physical Vacuum and Multi-Instanton Resummation
Before we formulate the dynamics of the Fermi string on the minimal surface, we must clarify the physical origin of this rigid geometric background. The MM (MM) loop equations serve as the exact quantum equations of motion for Planar QCD. Like any equations of motion, they possess an inherent ambiguity corresponding to the choice of the physical vacuum. The mathematical freedom to multiply the perturbative loop functional by a zero mode exactly reflects this physical vacuum ambiguity.
In the traditional gauge-field picture, the nonperturbative QCD vacuum is a highly complex mixed state, often modeled as a nontrivial distribution in the multi-instanton space. Because any single multi-instanton solution explicitly breaks translation invariance by localizing topological charge at specific spacetime points, restoring Poincaré symmetry requires an exceedingly complicated integration over the entire multi-instanton moduli space (positions, sizes, and color orientations).
The loop-space formulation achieves this summation geometrically. Just as instantons are characterized by the self-duality of the gauge field strength ($F_{\mu\nu}=\pm*F_{\mu\nu}$), our exact zero mode is characterized by the self-duality of the loop-space area derivative.
Thus, the complex mixed state of multi-instantons in coordinate space simplifies to a single, fully translation-invariant, exact geometric object in loop space: the Hodge-dual minimal surface.
Note that the mere Hodge duality condition leaves the freedom of an arbitrary topological charge and arbitrary moduli of the multi-instanton solution. Therefore, our Hodge-dual minimal surface, by satisfying this requirement, represents a general statistical mixture of multi-instanton solutions, implicitly summed over topological charges and integrated over moduli with some unknown measure. This implicit integration is precisely what provides the exact translation invariance and spatial isotropy of the minimal area solution. The profound advantage of the loop-space approach is that we do not need to know the explicit details of this distribution of topological charge and multi-instanton moduli: the loop equations can be solved directly without ever specifying this measure.
Therefore, the rigid background upon which our worldsheet fermions propagate is not merely a mathematical ansatz. It is the exact holographic representation of the translation-invariant, nonperturbative QCD vacuum, naturally incorporating the sum over topological sectors that is analytically prohibitive in the standard gauge-field path integral.

### 23 Mar 2026 v4

@@ Abstract
We present a local, asymptotically free solution of the planar MM loop equations in the continuum limit with full Lorentz invariance. The solution is constructed by quantizing internal Majorana fermions (referred to here as “elves”) on a rigid Hodge-dual minimal surface. These worldsheet degrees of freedom provide the mechanism to satisfy the unintegrated vector loop equations, with the Pauli principle enforcing planar factorization. In the local limit, the theory reduces to a confining analytic twistor-string representation. By analyzing the monodromy structure of the complexified effective action, we show that the discrete mass spectrum is organized by topological data associated with twistor singularities. The simplest sector with one branch point yields parametric Regge trajectories expressed by a universal formula with trigonometric functions. These trajectories are approximately linear for light mesons over a broad range and are in agreement with experimental data for 36 light, strange, charmed, and bottom mesons. The asymptotic behavior of the trajectory $J=\alpha(M^{2})$ and its subleading corrections arise from the interaction between the Liouville term and the twistor monodromy, without introducing additional assumptions about string excitations. In our solution, the QCD mass spectrum follows from a generalized eigenvalue problem in complexified phase space, reducing the problem to classical geometry. Within this framework, the large-$N_{c}$ Master Field is realized as a trajectory in twistor space.
@@ 1 Introduction
The identification of a string description of large-$N_{c}$ QCD has been a long-standing problem since the formulation of the MM loop equations [^1] The Makeenko–Migdal equations, referred to hereafter as the MM equations, exhibit factorization properties characteristic of a string description. The foundational derivation is due jointly to the author and Yuri Makeenko.. A continuum realization remains nontrivial. A technical challenge is the ultraviolet singular behavior of the coordinate-space Wilson loop functional $W[C]$. Even after renormalization, $W[C]$ exhibits cusp and perimeter divergences, and the loop equation contains a contact term $\delta^{(4)}(x-y)$ at self-intersections. In the coordinate representation, the loop equation is local, while perturbative solutions are nonlocal and develop logarithmic singularities at cusps and intersections.
In this work, we adopt a different approach. Rather than attempting to renormalize the coordinate-space Wilson loop as an observable, we formulate the dynamics in momentum loop space. The momentum loop functional $W[P]$, introduced in [26], is the functional Fourier transform of $W[C]$. In this representation, the planar loop equation becomes an algebraic–differential functional equation without coordinate-space delta functions. The geometric singularities of $W[C]$ are integrated out, and the continuum equations admit a local formulation appropriate for quark-loop amplitudes.
The aim of this second paper in the Geometric QCD series is to provide the dynamical construction complementing the geometric framework developed in Part I [32]. In that work, the Hodge-dual additive minimal surface $S_{\chi}[C]$ was shown to generate a confining factor $\exp\left(-\kappa S[C]\right)$ as a zero mode of the loop operator. Here, we introduce worldsheet degrees of freedom on this rigid surface and construct a functional $W[C]$ that satisfies the planar loop equations. Our framework combines four elements:
@@
    Fermi string (“Elfin theory”). Internal Majorana degrees of freedom on the worldsheet implement planar topology. The Pauli principle cancels nonplanar intersection configurations, leaving planar contributions.
@@
    Rigid Hodge-dual minimal surface. The fermions propagate on a surface determined by the boundary loop. There is no integration over worldsheet metrics, and the geometry is fixed by the loop data.
@@
    Momentum loop dynamics. In momentum loop space, the loop equation takes an algebraic–differential form. A representation in terms of the Cuntz algebra exists [26], but a one-dimensional Master Field construction exhibits a rank deficiency at order $\mathcal{W}^{(8)}$ in the Voiculescu expansion, indicating the need for additional degrees of freedom.
  4. 4.
    Twistor parametrization. The reduced loop-space measure is parametrized by boundary twistor variables. This yields a twistor parametrization in which a boundary sigma model is coupled to a Liouville field determined by analytic continuation from the boundary.
@@
Equation solving vs. model building. The starting point of this work is the planar MM loop equations, treated as defining equations of large-$N_{c}$ Yang–Mills theory. The goal is to identify variables and a functional measure in which these equations admit a finite continuum realization. The ingredients of the construction are constrained by this requirement. The Hodge-dual minimal surface appears because its area satisfies the loop equation, while the Majorana determinant satisfies planar factorization, reproducing planar graphs of QCD.
@@
Role of vector loop equations and the extrinsic surface geometry. A central technical point underlying this construction is that the full, unintegrated vector form of the MM loop equation contains essential information that is not captured by scalar or integrated formulations. In particular, when the momentum loop functional $W[P]$ is expanded in a one-dimensional Taylor–Magnus series, the resulting algebraic system becomes overconstrained at order $\mathcal{W}^{(8)}$. This is not a limitation of the expansion technique but a genuine inconsistency: the number of independent tensor constraints exceeds the number of available scalar degrees of freedom.
This obstruction indicates that a one-dimensional loop functional does not possess sufficient internal structure to represent a solution of the vector loop equations. The missing tensorial degrees of freedom are associated with the extrinsic geometry of the embedding surface. In the present work, these are provided by the boundary twistor variables and the fermionic determinant on the Hodge-dual minimal surface. From this perspective, the construction provides the geometric structure required for a consistent local solution of the planar loop equations in the continuum limit.
@@ 2 Zero modes of Loop equation
The theory of Hodge-dual minimal surfaces and its application to the MM loop equation is summarized here. For detailed discussion and proofs of the relevant theorems, see [32]; further details on the loop calculus without divergences are provided in [29, 33].
@@ 2.1 Self-dual area derivative and loop equation
As suggested in [32], we examine the self-dual minimal surface, which we will later treat as a background geometry for the QCD string with attached Fermi degrees of freedom. The area derivative for a functional $W[C]$ is defined as the discontinuity of the second variation of the area:
@@
This implements the geometric definition of [20, 23] and is formally equivalent to the functional derivative implementation in [36]. A significant feature of this definition is the absence of divergences and its parametric invariance. In previous implementations, the zero modes were often complicated by kinematic divergences.
The loop equation for the Yang–Mills gradient flow reduces to the loop space diffusion equation [29], admitting the solution $W=\exp\left(-\kappa S[C]\right)$ provided $S[C]$ is a zero mode:
@@
There exists a Hodge-dual solution analogous to the Yang–Mills multi-instanton. The loop space diffusion equation (2.3) is satisfied for any functional $S[C]$ with a self-dual (SD) or anti-dual (ASD) area derivative:
@@
The Bianchi identity (2.5) holds as a Jacobi identity for a triple commutator [29]. The exponential of the Hodge-dual minimal area serves as a factor in a general solution to the loop equation. This zero mode relies on the Leibniz property of the loop operator:
@@
This property leads to:
@@
We identify a self-dual minimal surface that is additive at self-intersections. This additivity ensures the confining factor $\exp\left(-\kappa S[C]\right)$ is compatible with the chain of loop equations.
2.2 Compatibility with the MM loop equations
The chain of loop equations for Wilson loops at finite $N_{c}$ [21, 23] utilizes the same loop operator (2.3) as the Yang–Mills gradient flow. We consider the leading order of the $1/N_{c}$ expansion using the loop calculus developed in [29]. The MM loop equation is structured as follows:
Figure 1: The MM equation, with the loop diffusion operator on the left side and the integral term on the right side. The double line represents the delta function $\delta^{4}(x-y)$ at self-intersections.
@@ 2.2 Compatibility with the MM loop equations
where $\lambda=N_{c}g_{0}^{2}$. This equation is understood in the sense of distributions. The confining factor $\exp\left(-\kappa S[C]\right)$ factors out by additivity and cancels on both sides of the equation independently of the splitting points $x,y\in C$.